Here is Question .....


x + y + z = 2

x - y + z = 1

x - y - z = 1


solve for x, y and z

Grade:

2 Answers

raj kaushik
15 years ago

This Question is quite simple ....i solved it ...

check out the solution

equation 1 + equation 2

=>       2(x+z) = 3
             x+z = 3/2
equation 1 +equation3

=>       2x = 3
=>       x = 3/2
equation 2+quation3

=>       2(x-y) =1
             x-y = 1/2
             y = 1
equationj 1 substitute the values of x and y

=>        z=-1/2

anirudh alameluvari
14 years ago

                                                                                     

                                                                                                 x+y+z = 2  ;   x-y+z=1 ;  x-y-z=1

                                                                                                          x+y+z+x-y+z=3 

                                                                                                             2(x+z) = 3

                                                                                                               x+z=3/2 

                                                                                                            x+y+z+x-y-z = 3

                                                                                                                    2x =3                  

                                                                                                                      x=3/2

                                                                                                                       3/2+z=3/2                                                              

                                                                                                                             z=0

                                                                                                                            x-y-z=1

                                                                                                                            3/2-y-0=1

                                                                                                                              y=1/2

                                                                                                                    (x,y,z) = (3/2,1/2,0)